Asiatic Express is a US firm and is considering starting a project in Malaysia. The project would end in 4 years. The expected (required) rate of return is 20%. It requires an initial investment of $100,000 Malaysian dollars and the following yearly cashflows (all in Malaysian dollars) are: $30,000 in year 1,

$50,000 in year 2, $70,000 in year 3, $90,000 in year 4.

Question 2
What is the cash flow to the parent in US dollars for year 1, year 2, and year 3? (3 points)

Question 3
What is the cumulated NPV at the end of year 3? (4 points)

(use 2 decimal places for answers where required, remember to convert all in US dollars)

Question 4
For Asiatic Express’s new project, in year 1, if the demand of the product is 1,500 units and price per unit is $25, what is the total revenue? (2 points)

Question 5
What are the Projected expenses + all taxes amount for Asiatic Express for year 1? (2 points)
(Hint: Revenue - cash flow)
Answer #1

2. One Malaysian Dollar = .23 US $

Thus for year 1 , Cash Flow in US $ will be - 30,000*.23 = 6,900$

For Year 2 , Cash Flow will be - 50,000 * .23 = 11,500$

For Year 3 , Cash Flow will be - 70,000 * .23 = 16,100$

3.- Initial Investment in US$ = 23,000$

- Cash Flow for year 1 = 6,900$

- Cah Flow for year 2 = 11,500$

- Cash Flow for year 3 = 16,100$

NPV after 3 year = casf flows/ (1+r)*i

Thus NPV after 3 year is 53.24$

4 - Demand in Year 1 is 1500 units

Price per unit = 25$

Total Revenue = 1500 * 25 = 37,500$

5 - Projected Expenses for year 1 ( Revenue of year 1 - Cash Flow of year 1)

= 37500$ - 30000$ = 7500$

Tax for year 1 = Revenue * Rate of return

= 37500 * 20% = 7500$

so, project expense + Total tax amount = 7500$ + 7500$ = 15,000$.